OpenStudy (anonymous):

If : 2 + 6 + 10 + 14 + 18 + 22 + 26 + 30 + 34 + 38 = 200, Then, Add any Five numbers and get 100..

OpenStudy (anonymous):

0+2+6+10+14+18+22+26+30+34+38=200 so 0+14+22+38+26=100 your question is solve ok

OpenStudy (anonymous):

(38 + 22) + (6 - 2)*(10) = 60 + 40 = 100

OpenStudy (anonymous):

0 is not there

OpenStudy (anonymous):

only plus

OpenStudy (nirmalnema):

6+10+14+38+22

OpenStudy (nirmalnema):

@saifpirjade is this correct

OpenStudy (anonymous):

nahi this add.. is 90 only worng ans

OpenStudy (anonymous):

nirmalnema plz send me ans

OpenStudy (anonymous):

@nirmalnema 6+10+14+38+22=90

OpenStudy (nirmalnema):

ok...sorry

OpenStudy (anonymous):

@nirmalnema tell friend in fb

OpenStudy (nirmalnema):

i m not your friend on FB

OpenStudy (nirmalnema):

can you use.....one no...twice

OpenStudy (anonymous):

no

OpenStudy (anonymous):

tell your friend on fb

OpenStudy (nirmalnema):

what...??

OpenStudy (anonymous):

fb per aapne friends ko ye ques. puch lo kisi ko jawb malum hoga to rly karega

OpenStudy (nirmalnema):

hm..

OpenStudy (anonymous):

plz..

OpenStudy (john_es):

\[2+4+38+26+30\]

OpenStudy (john_es):

Anohter \[4+14+18+26+38\]

OpenStudy (anonymous):

@John_ES but 4 is not there

OpenStudy (john_es):

My fault XD,

OpenStudy (anonymous):

send another ans

OpenStudy (sirm3d):

it's not possible

OpenStudy (john_es):

With these numbers is not possible, true.

OpenStudy (john_es):

Unless you can repeat some of them.

OpenStudy (anonymous):

can't you use one no twice

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

20+20+20+30+10=100